OpenStudy (anonymous):

I need help on these two questions since I don't quite understand. Someone explain the steps on figuring them out. "Find [fog][x] and [gof][x] 1. f(x)=2x-3 g(x)=x^2-2x 2. f(x)=x^2-9 g(x)=x+4

OpenStudy (anonymous):

first get rid of the circle, and write \[f\circ g(x)=f(g(x))\] then work from the inside out

OpenStudy (anonymous):

\[f(g(x))\] replace the general \(g(x)\) by the specific one you have, and write \[f(g(x))=f(x^2-2x)\]

OpenStudy (anonymous):

now comes the only tricky part \[f(x)=2x-3\] is like \(f(\diamondsuit)=2\diamondsuit -3\) replace \(\diamondsuit\) by \(x^2-2x\) using parentheses and get \[f(x^2-2x)=2(x^2-2x)-3\]

OpenStudy (anonymous):

then some algebra to clean it up \[2(x^2-2x)-3\] distribute the \(2\) to get \[2x^2-4x-3\]

OpenStudy (anonymous):

final answer is \[f\circ g(x)=2x^2-4x-3\]

OpenStudy (anonymous):

\[g\circ f(x)\] method is the same, algebra is a bit more \[g\circ f(x)=g(f(x))=g(2x-3)=(2x-3)^2-2(2x-3)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw hope you can now do the second one by mimicking the first two examples always works the same way